OpenStudy (anonymous):

Find the equation of the tangent line to f(x)=1/x^-2 at x=9

OpenStudy (anonymous):

can you clarify the equation?

OpenStudy (anonymous):

i mean 1 over square root of x

OpenStudy (freckles):

You need to find the derivative of 1/x first.

OpenStudy (anonymous):

i know and i forgot how to do it the long way

OpenStudy (freckles):

oh it is 1/sqrt(x)

OpenStudy (freckles):

you have to do it the long way ?

OpenStudy (anonymous):

lol i forgot also....

OpenStudy (anonymous):

Yes. WE havent learned the shortcut yet but i have

OpenStudy (freckles):

\[f'(x)=\lim_{h \rightarrow 0}\frac{f(x+h)-f(x)}{h}\]

OpenStudy (anonymous):

i tried that and i got stuck after plugging it in

OpenStudy (freckles):

\[f'(x)=\lim_{h \rightarrow 0}\frac{\frac{1}{\sqrt{x+h}}-\frac{1}{\sqrt{x}}}{h} =\lim_{h \rightarrow 0}\frac{1}{h} \cdot (\frac{1}{\sqrt{x+h}}-\frac{1}{\sqrt{x}})\]

OpenStudy (freckles):

First step : Combine fractions in the ( )

OpenStudy (freckles):

Second step: Combine like terms on top

OpenStudy (freckles):

Third step: Cancel a factor of h on bottom with the factor of h on top.

OpenStudy (freckles):

Last step: Replace remaining h's with 0.

OpenStudy (anonymous):

the combined fractions for the first step are sqrt x - sqrt x+h all over sqrtx * sqrt x+h right?

OpenStudy (freckles):

oh i missed step rationalize numerator

OpenStudy (freckles):

That should be first and half step lol

OpenStudy (anonymous):

how would i do that

OpenStudy (freckles):

\[f'(x)=\lim_{h \rightarrow 0}\frac{1}{h}(\frac{\sqrt{x}-\sqrt{x+h}}{\sqrt{x+h} \sqrt{x}}) \cdot \frac{\sqrt{x}+\sqrt{x+h}}{\sqrt{x}+\sqrt{x+h}}\] multiply by top's conjugate on top and bottom

OpenStudy (freckles):

do not distribute on bottom just on top

OpenStudy (anonymous):

what would remain after that?

OpenStudy (freckles):

multiply the top and see

OpenStudy (freckles):

recall (a-b)(a+b)=a^2-b^2

OpenStudy (anonymous):

x-x+h?

OpenStudy (freckles):

close x-(x+h)

OpenStudy (anonymous):

oh ok yeah I figured

OpenStudy (anonymous):

and then thats over the current denominator of sqrtx+h * sqrtx

OpenStudy (anonymous):

how would i fix up the bottom?

OpenStudy (freckles):

h/h=1

OpenStudy (anonymous):

what? how does h/h fix the sqrts on the bottom

OpenStudy (freckles):

why do the sqrt need to be fixed?

OpenStudy (anonymous):

so its (1/h) * (x-x+h)/ (sqrt x+h)(sqrtx)

OpenStudy (freckles):

again we have x-(x+h) on top you are also missing the other part of the bottom

OpenStudy (freckles):

x-x-h=-h and again h/h=1

OpenStudy (anonymous):

the bottom part of the congragate?

OpenStudy (freckles):

\[f'(x)=\lim_{h \rightarrow 0}\frac{1}{h}(\frac{\sqrt{x}-\sqrt{x+h}}{\sqrt{x+h} \sqrt{x}}) \cdot \frac{\sqrt{x}+\sqrt{x+h}}{\sqrt{x}+\sqrt{x+h}} \\ =\lim_{h \rightarrow 0}\frac{1}{h}\frac{x-(x+h)}{\sqrt{x}\sqrt{x+h}(\sqrt{x}+\sqrt{x+h})}\]

OpenStudy (anonymous):

yeah

OpenStudy (freckles):

x-x-h=-h h/h=1

OpenStudy (freckles):

You are on that third step I mentioned

OpenStudy (anonymous):

so the bottom cancels out?

OpenStudy (freckles):

:(

OpenStudy (freckles):

doesn't x-(x+h)=-h?

OpenStudy (anonymous):

yes

OpenStudy (freckles):

do you see a factor h on bottom?

OpenStudy (anonymous):

oh dear god

OpenStudy (anonymous):

idk what i was thinking

OpenStudy (anonymous):

ok yeah now i completely get it i guess i was trying to solve the problem and get a number rather than just find the derivative

OpenStudy (freckles):

\[\lim_{h \rightarrow 0}\frac{- \cancel{h}}{\cancel{h} \sqrt{x}\sqrt{x+h}(\sqrt{x}+\sqrt{x+h})}\]

OpenStudy (freckles):

now you can replace the h's with 0

OpenStudy (anonymous):

Ok well thank you and i apologize for my being dumb for about 20min there lol

OpenStudy (freckles):

It's okay. I was eating oreos part ways through it.

OpenStudy (anonymous):

is there a way i "friend" you for help i might need in the future?

OpenStudy (freckles):

friend you is a cool term There is fanning You can message me and see if i'm on when you need help

OpenStudy (freckles):

I will respond if I'm on

OpenStudy (anonymous):

Thank you